The CHL2 cryo is back on and cold! The problems on the cryo system from the power outage has been solved. Machine is being tuned with the expectation of beam delivery some time later this evening. Tonight's owl shift people should report for their shifts; swing shift people should report for half of their shifts, 8:00 pm - 12:00 am. Thanks, Mahbub - RC